# 4D-VLA: Spatiotemporal Vision-Language-Action Pretraining with Cross-Scene Calibration

### v1

@@ 2 Related works
Vision-language-action models The vision-language-action (VLA) model represents a significant advancement in vision-language research, enabling more complex and interactive tasks aimed at facilitating real-world environment interactions.  chi2023diffusionpolicy ; huang2023voxposer ; huang2024rekep complete tasks by directly predicting trajectories. wen2023any ; bharadhwaj2024track2act ; jiang2022vima ; li2023vision ; xu2024flow ; huang2023embodied predict the robot’s current actions to enable closed-loop control, while 3d-vla ; wu2023unleashing enhance action prediction by training a world model to forecast future states.
@@ 3.2 Spatial-aware visual tokens
In our method, the input image $\textbf{I}\in\mathbb{R}^{3\times h\times w}$ is first encoded by $\mathcal{E}$ into a feature map with a downsampling rate of $c$, yielding $\textbf{f}_{v}=\mathcal{E}(\textbf{I})\in\mathbb{R}^{k\times\frac{h}{c}\times\frac{w}{c}}$, where $k$ denotes the feature dimension. Next, we obtain a downsampled depth map $\mathbf{D}\in\mathbb{R}^{\frac{h}{c}\times\frac{w}{c}}$, which assigns depth values to each corresponding feature volume. Using the camera’s extrinsic $[\mathbf{R}|\mathbf{T}]$ and intrinsics $\mathbf{K}$, we back-project the depth value into 3D coordinates $\mathbf{P}_{w}\in\mathbb{R}^{3\times\frac{h}{c}\times\frac{w}{c}}$ within the world (or robot) coordinate system:
@@
where $\bm{e}^{text}$ represents the text instruction tokens. Each sampled frame $i$ contributes a token pair $[\bm{e}^{T}_{i}\mid\bm{e}^{ST}_{i}]$, with the temporal encoding token preceding the spatial vision token. This design ensures that the model effectively captures temporal relationships while maintaining spatial-awareness.
@@
Initialize $\mathcal{H}\leftarrow[t]$triangleright Start with the current frame
Initialize $\mathbf{S}\leftarrow[-\inf]$triangleright Similarity list
for $j=1$ to $n-1$ do
@@
Append $t-j$ to $\mathcal{H}$; Append $s$ to $\mathbf{S}$
@@
Append $t-j$ to $\mathcal{H}$; Append $s$ to $\mathbf{S}$
@@
Append $t-j$ to $\mathcal{H}$; Append $s^{\prime}$ to $\mathbf{S}$
@@ 3.4 Loss functions
To accelerate control policy generation, we use an action head with two MLP layers that predict the action $[\Delta{\hat{\bm{x}}},\Delta{\hat{\bm{\theta}}},{\hat{g}}]$ using the hidden features of the VLM Transformer’s last token.
@@ 4 Experiments
In this section, we introduce the dataset and simulation environment, followed by a detailed description of our pretraining and fine-tuning process. Our model is first pretrained on real-world datasets and then fine-tuned using both simulation and real-world data. Finally, we perform closed-loop evaluations across diverse environments, measuring task success rates to assess overall performance.
@@ 4.2 Pretraining setups
Data process. We pretrain our model on the DROID khazatsky2024droid dataset. RGB-D frames are resized to 448$\times$252, and each trajectory is uniformly downsampled to 100 actions. We remove frames with unchanged proprioception, specifically the stationary frames, and exclude trajectories with a total action count exceeding 600. Actions are defined as the difference between the end-effector’s current and target states, with translation scaled by 15 and rotation (Euler angles) by 5 for normalization.
@@ 4.3 LIBERO evaluation
Fine-tuning details. Unlike the pretraining phase, we used the simplest input settings to enable our model to learn the interaction effects between 3D information and historical data on actions. In the subsequent fine-tuning phase, we set the number of sequential frames $k=5$ with a window size $n=20$. The current state RGB-D image is resized to 448 $\times$448, while historical images are resized to 224$\times$224 to optimize memory usage and model efficiency. Following kim2024openvla , we normalize the ground truth action label using the 99th and 1st percentiles. We employed a cosine learning rate scheduler with a learning rate of 4e-5, using a batch size of 128 and training for 20 epochs. $\lambda_{d}$ is set to 0. During training, we activated all network parameters except the vision encoder.
@@
Angle $0^{\circ}$ $60^{\circ}$ $120^{\circ}$ $270^{\circ}$ $300^{\circ}$ $330^{\circ}$
@@ 5 Discussion
Building upon previous experiments, we further analyze three key questions: (i) The Role of historical information -How does historical context influence model’s effectiveness? (ii) Ablation study on model components -How do different architectural components contribute to the overall performance of the model? (iii) The impact of coordinate system chaos -How does coordinate system inconsistency affect the model’s performance, and can the introduction of a 3D representation mitigate this issue? The first two questions are addressed in Sec. 5.1 and Sec. 5.2, respectively, while the third is discussed in Appendix. 7.1. Our discussion begins with the simple model, i.e., InternVL-4B with an MLP action head, using single RGB image as the vision input.
@@ 5.1 Exploring historical information utilization
Table 5: Ablation results . Left: Ablation results on action head. Right: Ablation on 3D coordinate embedding, proprioceptive token and pretraining.
5.2 More ablation study
@@
We conduct additional ablation experiments on LIBERO-LONG. For the action heads (Tab. 5 left), the MLP head achieves the highest inference speed with a relatively high success rate. In contrast, the auto-regression approach, which uses text representations of actions and requires reasoning over multiple tokens, is slower.
@@
Results for the other proposed components and operations are presented in Tab. 5 right. From the table, proprioceptive tokens have a negative impact on the model, which may be due to the high correlation between proprioceptive states (e.g., joint positions, velocities) and action labels. This tight coupling can lead the model to overfit to dataset-specific motion patterns, rather than learning generalizable visuomotor representations.
6 Conclusion

### v2

@@ 2 Related works
Vision-language-action models The vision-language-action (VLA) model represents a significant advancement in vision-language research, enabling more complex and interactive tasks aimed at facilitating real-world environment interactions.  chi2023diffusionpolicy; huang2023voxposer; huang2024rekep complete tasks by directly predicting trajectories. wen2023any; bharadhwaj2024track2act; jiang2022vima; li2023vision; xu2024flow; huang2023embodied predict the robot’s current actions to enable closed-loop control, while 3d-vla; wu2023unleashing enhance action prediction by training a world model to forecast future states. Some works radosavovic2024humanoid; radosavovic2023robot also drives VLA policies using simple concatenation of historical observations.
@@ 3.2 Spatial-aware visual tokens
In our method, the input image $\textbf{I}\in\mathbb{R}^{3\times h\times w}$ is first encoded by $\mathcal{E}$ into a feature map with a downsampling rate of $c$, yielding $\textbf{f}_{v}=\mathcal{E}(\textbf{I})\in\mathbb{R}^{k\times\frac{h}{c}\times\frac{w}{c}}$, where $k$ denotes the feature dimension. Next, we obtain a downsampled depth map $\mathbf{D}\in\mathbb{R}^{\frac{h}{c}\times\frac{w}{c}}$, which assigns depth values to each corresponding feature volume. Using the camera’s extrinsic $\mathbf{R}|\mathbf{T}]$ and intrinsics $\mathbf{K}$, we back-project the depth value into 3D coordinates $\mathbf{P}_{w}\in\mathbb{R}^{3\times\frac{h}{c}\times\frac{w}{c}}$ within the world (or robot) coordinate system:
@@
where $\bm{e}^{text}$ represents the text instruction tokens. Each sampled frame $i$ contributes a token pair $\bm{e}^{T}_{i}\mid\bm{e}^{ST}_{i}]$, with the temporal encoding token preceding the spatial vision token. This design ensures that the model effectively captures temporal relationships while maintaining spatial-awareness.
@@
Initialize $\mathcal{H}\leftarrow[t]$ $\triangleright$ Start with the current frame
Initialize $\mathbf{S}\leftarrow[-\inf]$ $\triangleright$ Similarity list
for $j=1$ to n1n-1 do
@@
Append tjt-j to $\mathcal{H}$; Append $s$ to $\mathbf{S}$
@@
Append tjt-j to $\mathcal{H}$; Append $s$ to $\mathbf{S}$
@@
Append tjt-j to $\mathcal{H}$; Append $s^{\prime}$ to $\mathbf{S}$
@@ 3.4 Loss functions
To accelerate control policy generation, we use an action head with two MLP layers that predict the action $\Delta{\hat{\bm{x}}},\Delta{\hat{\bm{\theta}}},{\hat{g}}$] using the hidden features of the VLM Transformer’s last token.
@@ 4 Experiments
We first introduce the datasets and simulation environment, then describe pretraining and fine-tuning. Our model is pretrained on real-world data and fine-tuned with both simulation and real-world trajectories. We run closed-loop evaluations in diverse environments and report task performance.
In addition, we evaluate on the ARM4R niupre benchmark for a direct comparison with ARM4R. Note that ARM4R pretrains with 3D inputs, while our approach differs in supervision and architecture; results and a detailed discussion are provided in Appx. 7.1.
@@ 4.2 Pretraining setups
[…]
@@ 4.3 LIBERO evaluation
Fine-tuning details. Unlike the pretraining phase, we used the simplest input settings to enable our model to learn the interaction effects between 3D information and historical data on actions. In the subsequent fine-tuning phase, we set the number of sequential frames $k=5$ with a window size $n=20$. The current state RGB-D image is resized to 448 $\times$448, while historical images are resized to $\times224$ to optimize memory usage and model efficiency. Following kim2024openvla, we normalize the ground truth action label using the 99th and 1st percentiles. We employed a cosine learning rate scheduler with a learning rate of 4e-5, using a batch size of 128 and training for 20 epochs. $\lambda_{d}$ is set to 0. During training, we activated all network parameters except the vision encoder.
@@
Angle $0^{\circ}$ $60^{\circ}$ $120^{\circ}$ $270^{\circ}$ $00^{\circ} 33$$30^{\circ}$
@@ 4.5 Real-world evaluation
4.6 Multi-view real-world evaluation
In this section, we conduct additional real-world experiments under a multi-view camera setup. We design two more challenging tasks to evaluate the model’s generalization ability with respect to: (i) variations in object locations together with the changed background environments; (ii) inputs from novel camera viewpoints. We use 4 fixed cameras to capture each demonstration from different angles, collecting 50 trajectories per task per camera—resulting in a total of 200 trajectories per task for training. All models are trained for 20 epochs, and performance is measured by success rate.
Figure 5: Our multi-view real-world experiment settings. These settings aim to evaluate the model’s out-of-distribution and novel-view generalization ability.
Method In-View Cross-View Avg.
Angle $0^{\circ}$ $90^{\circ}$ $180^{\circ}$ $225^{\circ}$ $\Delta 15^{\circ}(-15^{\circ})$ $\Delta 25^{\circ}(-25^{\circ})$ $\Delta 45^{\circ}(135^{\circ})$
OpenVLA kim2024openvla 25 15 30 10 30 10 5 18
4D-VLA (Ours) 60 50 65 65 50 55 40 55
Table 4: Real-world multi-view evaluation. We test our model’s spatial generalization across varying viewpoints and object layouts. 4D-VLA shows strong in-view and cross-view performance, highlighting its robustness under real-world distribution shifts.
Task descriptions.
These two more challenging tasks are shown in Fig. 5. Task 1: Out-of-distribution generalization. The robot is tasked with placing a yellow cube into a plate under conditions where both the spatial configuration and the surrounding environment differ from those seen during training. These variations include changes in the plate’s position, the presence and location of distractor objects (e.g., other cubes). This task evaluates the model’s ability to generalize to unseen object arrangements and background contexts, testing its robustness in real-world deployments beyond the training distribution. Task 2: Novel-view generalization. Similar to Task 1, the robot is asked to place a yellow cube into a plate, with spatial setup and surroundings differing from training. However, during inference, data input is captured exclusively from an additional novel, unseen camera viewpoint that was not used during training. This task evaluates the model’s viewpoint robustness—its ability to generalize across camera perspectives and accurately interpret the scene from unfamiliar angles. Success in this task reflects strong spatial understanding and invariance to viewpoint changes, both critical for real-world multi-camera deployment. To simplify the setup, the target block is only moved within a small spatial range, while the background is fully randomized.
Evaluation metrics.
Each multi-view task is evaluated over 20 trials. In every trial, both the background and object positions are randomly shuffled to assess the model’s robustness and generalization. The evaluation metric is the task success rate, computed as the ratio of successful trials to the total number of trials.
Experiments results.
As shown in Tab. 4, 4D-VLA significantly outperforms OpenVLA in both in-view and cross-view settings, demonstrating strong generalization to viewpoint shifts and layout variations. In the in-view setting, where the camera is fixed but object layouts change, our model maintains consistently high success rates, indicating robustness to spatial perturbations. In the more challenging cross-view setting involving unseen viewpoints, 4D-VLA continues to perform stably across different angles. Although performance slightly drops at larger viewpoint shifts (e.g., $\Delta$45°), it remains stable compared to OpenVLA, whose success rate fluctuates more severely under such conditions. These results suggest that our model effectively captures spatial consistency across views, leading to more reliable visuomotor control in real-world environments.
@@ 5 Discussion
Building upon previous experiments, we further analyze three key questions: (i) The Role of historical information -How does historical context influence model’s effectiveness? (ii) Ablation study on model components -How do different architectural components contribute to the overall performance of the model? (iii) The impact of coordinate system chaos -How does coordinate system inconsistency affect the model’s performance, and can the introduction of a 3D representation mitigate this issue? The first question is addressed in Sec. 5.1; the second is covered in Appx. 8, and the third in Appx. 7.3. Our discussion begins with the simple model, i.e., InternVL-4B with an MLP action head, using single RGB image as the vision input.
@@ 5.1 Exploring historical information utilization
6 Conclusion
Acknowledgments
This work was supported in part by National Natural Science Foundation of China (Grant No. 62376060).
7 Appendix section
7.1 Comparison with ARM4R
Our approach differs from ARM4R in both design and practice. Instead of representation learning with full 3D point clouds and proprioception, we adopt an end-to-end, action-centric pretraining pipeline that maps vision–language inputs directly to low-level actions. Inputs are RGB with depth-aligned spatial coordinates (2D patches + $(x,y,z)$ embeddings), keeping the model lightweight and VLM-compatible while avoiding point-cloud encoders.
Method Meet-off-grill Push-buttons Place-wine Open-drawer Avg.
C2FARM-BC [james2022coarse] 20.0 72.0 18.0 20.0 32.5
LLARVA [niu2025llarva] 80.0 56.0 12.0 60.0 52.0
PerAct [shridhar2023perceiver] 84.0 48.0 12.0 80.0 56.0
ARM4R [niupre] 94.4 67.2 36.0 88.8 71.9
4D-VLA(Ours) 95.2 79.2 45.6 89.6 77.4
Table 6: RLBench results under the ARM4R protocol.
Analysis.
To balance efficiency and spatial coverage, we use our streaming memory bank sampling strategy that keeps 5 history frames within a 20-frame context. The current step uses dual views (front + wrist), while history stores only the front view. We also adopt OpenVLA’s random-crop augmentation. This configuration provides strong spatial grounding and good runtime, offering a strategy different from ARM4R yet comparably effective on its benchmark.
Note on supervision.
Our method leverages calibrated intrinsics/extrinsics during training and inference. While such calibration is standard in many robotic datasets, it is an additional source of supervision not explicitly used by ARM4R.
7.2 More ablation study
@@
Table 7: Ablations on heads and inputs (Libero-Long). Left: action head vs. FPS and success (MLP, autoregressive, diffusion). Right: effect of pretraining, 3D coordinate embedding, and proprioceptive tokens on success.
Sampling Method Success $\uparrow$ Latency (ms) $\downarrow$ Peak Mem (MB) $\downarrow$
Single frame 0.738 76.5 8342.5
Adaptive Pooling [shenlongvu] 0.604 150.7 8949.1
Grid Pooling [li2024llama] 0.620 208.8 8852.7
Q-Former (per-frame) 0.556 223.3 8812.8
MBS (Ours) 0.866 160.0 8682.9
Table 8: Frame sampling ablations on Libero-Spatial. MBS attains the highest success (0.866) with competitive efficiency, while single-frame is fastest and most memory-light but less accurate.
Action heads.
@@
We conduct additional ablations on LIBERO-LONG. For the action heads (Tab. 7, left), the MLP head yields the highest inference speed with a relatively high success rate. In contrast, the autoregressive head—predicting text tokens for actions—runs slower due to multi-token reasoning.
Pretrain and spatial components.
@@
Results for other components/operations are shown in Tab. 7 (right). Proprioceptive tokens hurt performance, likely because proprioceptive states (e.g., joint positions, velocities) are highly correlated with action labels, which encourages overfitting to dataset-specific motion patterns rather than learning generalizable visuomotor features.
Memory bank sampling.
We compare our memory bank sampling with three video-style samplers: Adaptive Pooling [shenlongvu], Grid Pooling [li2024llama], and a per frame Q-Former, using InternVL2.5-4B on Libero-Spatial (20 epochs). We report success, latency, and peak memory in Tab. 8. Adaptive Pooling caches the full trajectory and uses two thresholds (0.99 for frame selection and 0.96 for spatial compression). We omit its second-stage cross-modal filter due to low language diversity (10 tasks) and observe hyperparameter sensitivity and variable-length tokens that increase memory and reduce accuracy. Grid Pooling replaces language tokens with 8 learnable queries and compresses each frame to a $2{\times}2$ grid (4 tokens), yielding $20{\times}(4{+}8)$ historical tokens; it is simple but nonadaptive and can drop key spatial cues. Q-Former (per frame) uses 10 learnable queries per frame to extract features but converges slowly and discards too much spatial context, which hurts success. MBS (ours) operates online with a window of 20 and a memory of 5. It selects informative key frames from the stream, stores only 5 history frames at $224{\times}224$, and keeps the current frame at $448{\times}448$. As a result, MBS achieves the highest success (0.866) with competitive cost (160.0 ms and 8,682.9 MB). Single frame is faster (76.5 ms) but weaker (0.738). The other baselines are slower and less accurate. MBS fits VLA’s streaming and causal nature because it avoids full-clip caching, uses relative temporal encoding, and remains compatible with closed-loop control. It yields stronger long horizon performance without extra retraining or heavy compute.
7.3 Coordinate system chaos impact analysis
To assess the impact of coordinate system chaos on VLA’s performance, we conduct a controlled experiment. The experiment consists of two main steps: First, we deliberately introduce controlled chaos into the LIBERO environment. Then, we compare the performance of models with and without 3D information. Our findings demonstrate that chaotic coordinate transformations significantly degrade model performance while incorporating 3D information effectively alleviates it.
Figure 7: Success rates under varying coordinate chaos levels.
Chaos generation. To simulate the diverse viewpoints in the pretraining dataset—where the robot’s body is absent from the image, and its coordinate system varies unpredictably—we select trajectories from a specific task in LIBERO-SPATIAL and render each trajectory from 30 distinct viewpoints. For each trajectory and its corresponding viewpoint, we introduce coordinate system chaos by applying a random translation $\mathbf{t}\in\mathbb{R}^{3}$ and rotation $\mathbf{q}\in\mathbb{SO}(3)$ to the robot’s coordinate frame.
After applying the coordinate transformation, the gripper’s grasping state in the ground-truth action remains unchanged. However, the rotation offset $\Delta\bm{\theta}$ and translation $\Delta\bm{x}$, along with the proprioceptive and camera pose information, undergo the following transformation:
$\Delta\bm{\theta}^{\prime}$ $=\psi^{-1}(\mathbf{q}\psi(\Delta\bm{\theta})\mathbf{q}^{\top}),$ $\Delta\bm{x}^{\prime}$ $=\mathbf{q}\Delta\bm{x},$
$\bm{\theta}^{\prime}$ $=\psi^{-1}(\mathbf{q}\psi(\bm{\theta})),$ $\bm{x}^{\prime}$ $=\mathbf{q}\bm{x}+\mathbf{t},$
$\mathbf{R}^{\prime}$ $=\mathbf{q}\mathbf{R},$ $\mathbf{T}^{\prime}$ $=\mathbf{q}\mathbf{T}+\mathbf{t}.$ (5)
Here, the function $\psi:\mathbb{R}^{3}\to\mathbb{SO}(3)$ maps an Euler angle to its corresponding rotation matrix. The terms $\Delta\bm{\theta}^{\prime}$ and $\Delta\bm{x}^{\prime}$ denote the transformed action values, while $\bm{\theta}^{\prime}$, $\bm{\theta}$ and $\bm{x}^{\prime}$, $\bm{x}$ represent the transformed and original rotation and position, respectively. Additionally, $\mathbf{R}^{\prime}|\mathbf{T}^{\prime}]$ and $\mathbf{R}|\mathbf{T}]$ correspond to the transformed and original camera poses.
In the subsequent training process, we utilize the transformed action values $\Delta\bm{\theta}^{\prime}$ and $\Delta\bm{x}^{\prime}$, along with the transformed camera parameters, for model training.
Implementation details. We employ a simple baseline, controlling for the involvement of 3D information. The baseline model extracts tokens from a single RGB view, while the alternative model converts an RGB-D frame into spatial vision tokens as input for the LLM Transformer. During testing, we do not apply random rotations or translations to the world coordinate system.
Experimental results. We control chaos levels by adjusting the magnitude of random rotations. Level 0 applies no rotation, while levels 1–3 introduce random z-axis rotations of 15°, 30°, and 90°, respectively. Translation $\mathbf{t}$ is randomly set within a range of 0.5. As shown in Fig. 7, without chaos, both models perform well, with 3D information further boosting success rates. Notably, the 3D model shows lower variance across viewpoints. As chaos increases, the non-3D model’s performance drops sharply, while the 3D model maintains relatively high success—highlighting the value of 3D cues in handling coordinate system chaos.